Find the product of three consecutive even integers whose sum is 42?
Answer by algebrahouse.com(1659)

x = first integer
x + 2 = second consecutive even integer
x + 4 = third consecutive even integer {consecutive even integers increase by 2}

x + x + 2 + x + 4 = 42 {their sum is 42}
3x + 6 = 42 {combined like terms}
3x = 36 {subtracted 6 from each side}
x = 12 {divided each side by 3}
x + 2 = 14 {substituted 12, in for x, into x + 2}
x + 4 = 16 {substituted 12, in for x, into x + 4}

the three integers are 12, 14, and 16

Their product
= 12 x 14 x 16 {product means multiply}
= 2688 {multiplied}

For more help from me, visit: www.algebrahouse.com